# Gotta Hear Them All: Sound Source Aware Vision to Audio Generation

### v3

@@ Abstract
Vision-to-audio (V2A) synthesis has broad applications in multimedia. Recent advancements of V2A methods have made it possible to generate relevant audios from inputs of videos or still images. However, the immersiveness and expressiveness of the generation are limited. One possible problem is that existing methods solely rely on the global scene and overlook details of local sounding objects (i.e., sound sources). To address this issue, we propose a Sound Source-Aware V2A (SSV2A) generator. SSV2A is able to locally perceive multimodal sound sources from a scene with visual detection and cross-modality translation. It then contrastively learns a Cross-Modal Sound Source (CMSS) Manifold to semantically disambiguate each source. Finally, we attentively mix their CMSS semantics into a rich audio representation, from which a pretrained audio generator outputs the sound. To model the CMSS manifold, we curate a novel single-sound-source visual-audio dataset VGGS3 from VGGSound. We also design a Sound Source Matching Score to measure localized audio relevance. By addressing V2A generation at the sound-source level, SSV2A surpasses state-of-the-art methods in both generation fidelity and relevance as evidenced by extensive experiments. We further demonstrate SSV2A’s ability to achieve intuitive V2A control by compositing vision, text, and audio conditions. Our generation can be tried and heard at https://ssv2a.github.io/SSV2A-demo.
Figure 1: Our SSV2A perceives multimodal sound sources in a scene for V2A immersiveness and expressiveness.
Figure 2: Pipeline of SSV2A. We perceive sound sources prompted by vision, text, or audio and disambiguate them in the semantically learned Cross-Modal Sound Source Manifold, after which we mix them to generate an audio clip with immersiveness and expressiveness.
1 Introduction
@@
As multimedia consumption surges, generating sound for silent videos or still images attracts high demands in various industries [71]. The synthesized audio can complement a virtual reality scene [28], create Foley for films and games [11], and enrich single-modality visual contents for people with visual impairment [73]. By learning from the widespread visual-audio pairs in video data, recent methods can generate visually relevant audio clips for this vision-to-audio (V2A) task. However, most existing methods [24, 53, 12, 65, 37, 60, 45, 69, 61] only model the mapping between the global visual scene and audio output while overlooking local details.
@@
In reality, sound is produced and recognized from sounding objects, i.e., sound sources, locally present in a soundscape [38]. For instance, in a street the sound comes from individual vehicles and passengers as illustrated in Fig. 1. Humans also perceive audio immersiveness and expressiveness from these sound sources in a visual scene [16]. In practice, audio engineers leverage individual sound sources to intuitively control sound synthesis [49].
@@
Can a V2A synthesizer utilize sound source-aware conditions to obtain better generation quality and control? To answer this question, we present a Sound Source-Aware V2A (SSV2A) generator. We model our system in semantic spaces for learning efficiency and include multimodal conditions from text and audio to boost sound source controllability. As depicted in Fig. 1, the perception can also come from audio sound source as a loudspeaker and text source as “street ambient”. We present SSV2A’s pipeline in Fig. 2. SSV2A first perceives multimodal sound source conditions as CLIP [46] or CLAP [13] semantic embeddings with visual detection and cross-modal translation. We then project them to a Cross-Modal Sound Source (CMSS) Manifold to disambiguate each source semantic. By disambiguation, we require the CMSS manifold to (1) contrast the source semantics and (2) respect the audio characteristics of each sound source. After querying CMSS embeddings of individual sound sources, SSV2A learns an attention-based Sound Source Remixer to mix them into a CLAP audio embedding with rich sound source information. This representation is passed to a pretrained audio generator, AudioLDM [34], to synthesize the output audio waveform.
@@
As the CMSS manifold contrastively learns from single-sound-source visual-audio pairs to disambiguate sound source semantics, we filter the VGGSound [3] data with visual detection to form a novel dataset, VGGSound Single Source (VGGS3), that contains 106K high-quality single-sound-source visual-audio pairs. We also apply a novel Cross-Modal Contrastive Mask Regularization (CCMR) during manifold learning to retain rich CLIP-CLAP semantics by reducing CMSS contrastive influence on similar visual-audio sources with CLIP and CLAP priors. To effectively evaluate generation relevance, we introduce a Sound Source Matching Score (SSMS) to compute the F1 score of overlapping sound source labels on ground-truth and generated samples with an audio classifier.
@@
Both objective and subjective results show that SSV2A outperforms state-of-the-art methods in V2A synthesis. We also demonstrate SSV2A’s intuitive generation control in experiments by flexibly compositing multimodal sound source prompts from vision, text, and audio.
@@
  • We present a novel V2A generation framework, SSV2A. To the best of our knowledge, this is the first work to address V2A synthesis at the sound-source level. Extensive experiments show that SSV2A achieves state-of-the-art results in both V2A generation fidelity and relevance.
  • We explore how sound-source disambiguation can enhance SSV2A synthesis with the CMSS manifold, along with a novel CCMR mechanism to guide cross-modal contrastive learning with foundation model priors.
  • During manifold training, we curate a high-quality single-sound-source visual-audio dataset, VGGS3, which supports learning of various sound-source-related tasks.
  • In evaluating audio generation relevance, we introduce an effective SSMS metric aware of localized sound sources.
@@
2 Related Works
2.1 Vision-to-Audio Generation
Early V2A methods [44, 6, 73, 20, 5, 7] train a source-specific V2A model on each audio class and cannot generalize to open-domain V2A synthesis. As a precursor, recent SpecVQGAN [24] learns a discrete neural codec [58, 14] of source-agnostic audio features and autoregressively generates audio codes with a Transformer [59]. Following SpecVQGAN, Im2Wav [53] further details its audio codec into low-level and high-level features. MaskVAT [45] leverages a pretrained codec DAC [29] and predicts audio tokens with a Masked Generative Transformer [2]. Another line of methods employ Diffusion [23] models. CLIPSonic-IQ [12] queries CLIP [46] to condition its Diffusion process. Diff-Foley [37] contrastively learns a temporally-aligned visual-audio prior to guide video-audio synchronization. Draw-an-Audio [66] leverages loudness signal, text caption, and masked video conditions simultaneously. More recently, some methods bridge visual conditions to the prior of a pretrained audio generator for efficient V2A learning. V2A-Mapper [60] maps CLIP embeddings to CLAP [13] space, from which a pretrained AudioLDM [34] model synthesizes the audio signal. V2A-SceneDetector [67] extends V2A-Mapper to multi-scene video with a detection module. Seeing and Hearing [65] aligns ImageBind [17] visual embeddings to AudioLDM. FoleyCrafter [69] devises a timestamp predictor to enhance synchronization during bridging. Very recently, FRIEREN [61] and MMAudio [9] explore V2A generation with Rectified Flow Matching [36]. Most existing methods condition on global visual scenes for V2A synthesis. Some recent works [32] [31] leverage pixel-level conditions for V2A synthesis, partially describing visual sounding objects. In reality, human perceive object-level sound sources across modalities and time [38]. Such a sound source-aware V2A generator remains uninvestigated.
2.2 Contrastive Cross-Modal Alignment
Contrastive representation learning [18] has significantly advanced cross-modal representation alignment. CLIP [46] aligns text and image modalities by learning from abundant text-image pairs. Many aforementioned V2A methods [53, 45, 12, 69, 60] benefit from its semantically rich visual representations. Similarly, CLAP [13] learns from text-audio pairs and is used extensively in V2A generation [37, 65, 60, 66, 34, 35]. Aside from modality alignment, Diff-Foley [37] shows that it is possible to respect temporal alignment in the contrastive visual-audio representation to benefit video-audio synchronization. However, the entanglement of temporal features in this representation limits Diff-Foley in generalizing to image-to-audio synthesis. In this work, we focus on taming a contrastive representation for sound source disambiguation and leave the temporal alignment to a downstream temporal aggregation module.
@@ 2.2 Contrastive Cross-Modal Alignment
Figure 3: Detailed Schematics of SSV2A Modules. (a) We learn two projectors to map the CLIP-CLAP embeddings of single-source visual-audio pairs to a joint semantic space with contrastive guidance, forming our CMSS manifold. An auxiliary CLAP reconstruction encodes audio semantics into this manifold. (b) The Sound Source Remixer attends to the CMSS embeddings concatenated with their CLIP semantics, generating a single CLAP audio representation which is passed to AudioLDM. (c) We reuse the CMSS reconstructor to generate source-wise “track semantics” in CLAP space and refine the Remixer samples iteratively. (d) We train an additional Temporal Aggregation (TA) module to attend to positionally embedded SSV2A generations across video frames and enhance visual-audio synchronization.
3 Method
@@
Approximating an audio distribution $Q(\mathbf{A}|\mathbf{a})$, the audio generator AudioLDM [34] generates audio signals $\mathbf{A}$ from CLAP [13] audio semantics $\mathbf{a}$. For learning efficiency, we employ a pretrained $Q$ and synthesize $\mathbf{a}$ instead of $\mathbf{A}$. Conditioned on multimodal sound sources, our objective is to learn a conditional distribution:
@@
where $\left\{\mathbf{s}^{\text{vis}}_{i}\right\}$, $\left\{\mathbf{s}^{\text{text}}_{j}\right\}$, and $\left\{\mathbf{s}^{\text{aud}}_{k}\right\}$ denote respectively the semantic embedding sets of $I$ visual sound sources, $J$ text sources and $K$ audio sources encoded with CLIP [46] or CLAP. We term the acquisition of these semantics as Sound Source Perception in Fig. 2 (a) and discuss it in Sec. 3.1.
The most straightforward way to approximate Eq. 1 is to train a standalone model that maps the perceived CLIP-CLAP semantics directly to $\mathbf{a}$. However, two CLIP features ambiguate this direct learning: (1) the CLIP image space models global visual context rather than contrasting individual objects, and (2) CLIP learns only from text-image data, which lacks awareness of the sources’ audio traits. As an efficient solution, we learn a Cross-Modal Sound Source (CMSS) manifold as illustrated in Fig. 2 (b) to project the CLIP-CLAP embeddings to a joint semantic space where the local sound sources are disambiguated. We elaborate on this core stage of SSV2A in Sec. 3.2.
Finally, we attentively mix the CMSS embeddings together in Fig. 2 (c) to generate $\mathbf{a}$. This stage involves a Sound Source Remixer which we explain in Sec. 3.3.
3.1 Sound Source Perception
@@
3.2 Cross-Modal Sound Source Manifold
We contrastively learn the CMSS manifold from single-sound-source visual-audio pairs to project the perceived sound source semantics in Sec. 3.1 to a joint semantic space for disambiguation, as shown in Fig. 3 (a). The CMSS manifold naturally accommodates the multimodality of our perceptions due to the bridging of CLIP and CLAP.
@@ 3.2 Cross-Modal Sound Source Manifold
given a single-source visual-audio pair as $\left(\mathbf{V},\mathbf{A}\right)$ and its CLIP-CLAP embeddings as $\left(\mathbf{v},\mathbf{a}\right)$. $\mathbf{e}$ denotes the CMSS embedding. The projectors optimize a contrastive loss to attract visual-audio embeddings from the same sound-source pair and repel those from different sources. Following the symmetric contrastive guidance of CLAP [13], our main objective can be formulated for a batch of $N$ pairs as:
@@
where $\lambda_{1}$ is a weight hyperparameter and $\mathcal{L}_{kl}$ is the summed K-L losses. During training, we model all three modules $\upsilon\left(\cdot\right)$, $\phi\left(\cdot\right)$, and $\chi\left(\cdot\right)$ with residually connected MLPs and alternatively optimize the projectors and generator on each batch with $\mathcal{L}_{\text{fold}}$.
@@
3.3 Sound Source Remixer
@@
We model $\psi\left(\cdot\right)$ with a stack of self-attention layers and learn it from visual-audio pairs in VGGSound. The visual sources are perceived from each video’s central frame following Sec. 3.1. Each token sequence $\left\{\mathbf{x}_{m}\right\}$ is zero padded to a fixed length of $M=64$. To enhance generation diversity, a Classifier-free Guidance [22] is applied during training by randomly zeroing out tokens. We replace the classic attention with Efficient Attention [54] and detail this architecture in Sec. A.2.2. During inference, we set $\mathbf{v}=\mathbf{0}$ for sound source conditions from audio modality.

### v4

# Gotta Hear Them All: Towards Sound Source Aware Audio Generation

@@ Abstract
Audio synthesis has broad applications in multimedia. Recent advancements have made it possible to generate relevant audios from inputs describing an audio scene, such as images or texts. However, the immersiveness and expressiveness of the generation are limited. One possible problem is that existing methods solely rely on the global scene and overlook details of local sounding objects (i.e., sound sources). To address this issue, we propose a Sound Source-Aware Audio (SS2A) generator. SS2A is able to locally perceive multimodal sound sources from a scene with visual detection and cross-modality translation. It then contrastively learns a Cross-Modal Sound Source (CMSS) Manifold to semantically disambiguate each source. Finally, we attentively mix their CMSS semantics into a rich audio representation, from which a pretrained audio generator outputs the sound. To model the CMSS manifold, we curate a novel single-sound-source visual-audio dataset VGGS3 from VGGSound. We also design a Sound Source Matching Score to clearly measure localized audio relevance. With the effectiveness of explicit sound source modeling, SS2A achieves state-of-the-art performance in extensive image-to-audio tasks. We also qualitatively demonstrate SS2A’s ability to achieve intuitive synthesis control by compositing vision, text, and audio conditions. Furthermore, we show that our sound source modeling can achieve competitive video-to-audio performance with a straightforward temporal aggregation mechanism.
Demo Websitehttps://SSV2A.github.io/SSV2A-demo/
Introduction
@@
As multimedia consumption surges, generating sound for a silent scene attracts high demands in various industries (Zhao, Xia, and Togneri 2019). The synthesized audio can complement a virtual reality scene (Kern and Ellermeier 2020), create Foley for films and games (Di Donato and McGregor 2024), and sonify visual contents for people with visual impairment (Zhou et al. 2018). By learning from text-audio or visual-audio pairs, recent methods can generate highly relevant audio clips given conditions as texts, images or videos. However, most existing methods (Božić and Horvat 2024) only model the mapping between global visual scene and sound while overlooking local details.
Figure 1: Our SS2A perceives multimodal sound sources in a scene for V2A immersiveness and expressiveness.
@@
In reality, sound is produced and recognized from sounding objects, i.e., sound sources, locally present in a soundscape (McAdams 1993). For instance, in a street the sound comes from individual vehicles and passengers as illustrated in Fig. 1. Humans also perceive audio immersiveness and expressiveness from sound source interactions (Gaver 1993). In practice, audio engineers leverage sound sources to intuitively synthesize sounds (Russ 2012).
Figure 2: Pipeline of SS2A. We perceive sound sources prompted by vision, text, or audio and disambiguate them in the semantically learned CMSS Manifold, which are then mixed to generate an audio clip with immersiveness and expressiveness.
@@
Can an audio synthesizer utilize sound source-aware conditions to obtain better generation quality and control? To answer this question, we present a Sound Source-Aware Audio (SS2A) generator. As image offers sound sources with straightforward curation and composition, we choose it as the primary modality to condition SS2A. We model our system in semantic spaces for learning efficiency and include multimodal conditions from text and audio to boost sound source control. As depicted in Fig. 1, the perception can also come from audio sound source as a loudspeaker and text source as “street ambient”. We present SS2A’s pipeline in Fig. 2. SS2A first perceives multimodal sound source conditions as CLIP (Radford et al. 2021) or CLAP (Elizalde et al. 2023) semantic embeddings with visual detection and cross-modal translation. We then project them to a Cross-Modal Sound Source (CMSS) Manifold to disambiguate each source. By disambiguation, we require the CMSS manifold to (1) contrast the source semantics and (2) respect the audio characteristics of each sound source. After querying CMSS embeddings of individual sound sources, SS2A learns an attention-based Sound Source Remixer to mix them into a CLAP audio embedding with rich sound source information. This representation is passed to a pretrained audio generator, AudioLDM (Liu et al. 2023), to synthesize the output audio waveform.
@@
As the CMSS manifold contrastively learns from single-sound-source image-audio pairs to disambiguate sound source semantics, we filter the VGGSound (Chen et al. 2020a) data with visual detection to form a novel dataset, VGGSound Single Source (VGGS3), that contains 106K high-quality single-sound-source image-audio pairs. We also apply a novel Cross-Modal Contrastive Mask Regularization (CCMR) during manifold learning to retain rich CLIP-CLAP semantics by reducing CMSS contrastive influence on similar visual-audio sources with CLIP and CLAP priors. To effectively evaluate generation relevance, we introduce a Sound Source Matching Score (SSMS) to compute the F1 score of overlapping sound source labels on ground-truth and generated samples with an audio classifier.
@@
Both objective and subjective results show that SS2A achieves state-of-the-art performance in image-to-audio synthesis, indicating the benefits of sound source modeling. We demonstrate SS2A’s intuitive generation control by flexibly compositing multimodal sound source prompts from vision, text, and audio to synthesize immersive qualitative samples. We further showcase that our sound source modeling can be straightforwardly extended to competitive video-to-audio synthesis with a temporal aggregation mechanism.
@@
  • We present a novel framework, SS2A, addressing audio synthesis at the sound-source level. Extensive experiments show that our multimodal sound source modeling leads to state-of-the-art results in image-to-audio generation and competitive video-to-audio performance.
  • We explore how sound-source disambiguation can enhance SS2A synthesis with the CMSS manifold, along with a novel CCMR mechanism to guide cross-modal contrastive learning with foundation model priors.
  • During manifold training, we curate a high-quality single-sound-source image-audio dataset, VGGS3.
  • In evaluating relevance between generated and ground-truth audio signals, we introduce a novel SSMS metric to explicitly match their localized sound sources, proposing a new objective for fine-grained audio generation.
@@
Figure 3: Detailed Schematics of SS2A Modules. (a) We learn two projectors to map the CLIP-CLAP embeddings of single-source visual-audio pairs to a joint semantic space with contrastive guidance, forming our CMSS manifold. An auxiliary CLAP reconstruction encodes audio semantics into this manifold. (b) The Sound Source Remixer attends to the CMSS embeddings concatenated with their CLIP semantics, generating a single CLAP audio representation which is passed to AudioLDM. (c) We reuse the CMSS reconstructor to generate source-wise “track semantics” in CLAP space and refine the Remixer samples iteratively. (d) We train an additional Temporal Aggregation (TA) module to attend to positionally embedded SS2A generations across video frames and enhance visual-audio synchronization.
Related Works
Vision-to-Audio Generation
Early V2A methods (Owens et al. 2016; Chen et al. 2017; Zhou et al. 2018; Hao, Zhang, and Guan 2018; Chen et al. 2018, 2020b) train a source-specific V2A model on each audio class and cannot generalize to open-domain V2A synthesis. As a precursor, recent SpecVQGAN (Iashin and Rahtu 2021) learns a discrete neural codec (Van Den Oord, Vinyals et al. 2017; Esser, Rombach, and Ommer 2021) of source-agnostic audio features and autoregressively generates audio codes with a Transformer (Vaswani et al. 2017). Following SpecVQGAN, Im2Wav (Sheffer and Adi 2023) further details its audio codec into low-level and high-level features. MaskVAT (Pascual et al. 2024) leverages a pretrained codec DAC (Kumar et al. 2023) and predicts audio tokens with a Masked Generative Transformer (Chang et al. 2022). Another line of methods employ Diffusion (Ho, Jain, and Abbeel 2020) models. CLIPSonic-IQ (Dong et al. 2023) queries CLIP (Radford et al. 2021) to condition its Diffusion process. Diff-Foley (Luo et al. 2023) contrastively learns a temporally-aligned visual-audio prior to guide video-audio synchronization. Draw-an-Audio (Yang et al. 2024) leverages loudness signal, text caption, and masked video conditions simultaneously. More recently, some methods bridge visual conditions to the prior of a pretrained audio generator for efficient V2A learning. V2A-Mapper (Wang et al. 2024a) maps CLIP embeddings to CLAP (Elizalde et al. 2023) space, from which a pretrained AudioLDM (Liu et al. 2023) model synthesizes the audio signal. V2A-SceneDetector (Yi and Li 2024) extends V2A-Mapper to multi-scene video with a detection module. Seeing and Hearing (Xing et al. 2024) aligns ImageBind (Girdhar et al. 2023) visual embeddings to AudioLDM. FoleyCrafter (Zhang et al. 2024b) devises a timestamp predictor to enhance synchronization during bridging. Very recently, FRIEREN (Wang et al. 2024b) and MMAudio (Cheng et al. 2025) explore V2A generation with Rectified Flow Matching (Liu, Gong, and Liu 2023). MultiFoley (Chen et al. 2025) employs a diffusion transformer to jointly map multimodal conditions to audio. Most existing methods condition on global visual scenes for V2A synthesis. Some recent works (Li, Zhao, and Yuan 2024) (Li et al. 2024) leverage pixel-level conditions for V2A synthesis, partially describing visual sounding objects. In reality, human perceive object-level sound sources across modalities and time (McAdams 1993). Such a sound source-aware V2A generator remains uninvestigated.
Contrastive Cross-Modal Alignment
Contrastive representation learning (Hadsell, Chopra, and LeCun 2006) has significantly advanced cross-modal representation alignment. CLIP (Radford et al. 2021) aligns text and image modalities by learning from abundant text-image pairs. Many aforementioned V2A methods (Sheffer and Adi 2023; Pascual et al. 2024; Dong et al. 2023; Zhang et al. 2024b; Wang et al. 2024a) benefit from its semantically rich visual representations. Similarly, CLAP (Elizalde et al. 2023) learns from text-audio pairs and is used extensively in V2A generation (Luo et al. 2023; Xing et al. 2024; Wang et al. 2024a; Yang et al. 2024; Liu et al. 2023, 2024). Aside from modality alignment, Diff-Foley (Luo et al. 2023) shows that it is possible to respect temporal alignment in the contrastive visual-audio representation to benefit video-audio synchronization. However, the entanglement of temporal features in this representation limits Diff-Foley in generalizing to image-to-audio synthesis. In this work, we focus on taming a contrastive representation for sound source disambiguation and leave the temporal alignment to a downstream temporal aggregation module.
Method
@@
Approximating an audio distribution $Q(\mathbf{A}|\mathbf{a})$, the audio generator AudioLDM (Liu et al. 2023) generates audio signals $\mathbf{A}$ from CLAP (Elizalde et al. 2023) audio semantics $\mathbf{a}$. For learning efficiency, we employ a pretrained $Q$ and synthesize $\mathbf{a}$ instead of $\mathbf{A}$. Conditioned on multimodal sound sources, our objective is to learn a conditional distribution:
@@
where $\left\{\mathbf{s}^{\text{vis}}_{i}\right\}$, $\left\{\mathbf{s}^{\text{text}}_{j}\right\}$, and $\left\{\mathbf{s}^{\text{aud}}_{k}\right\}$ denote respectively the semantic embedding sets of $I$ visual sound sources, $J$ text sources and $K$ audio sources encoded with CLIP (Radford et al. 2021) or CLAP. We term the acquisition of these semantics as Sound Source Perception in Fig. 2 (a).
The most straightforward way to approximate Eq. 1 is to train a standalone model that maps the perceived CLIP-CLAP semantics directly to $\mathbf{a}$. However, two CLIP features ambiguate this direct learning: (1) the CLIP image space models global visual context rather than contrasting individual objects, and (2) CLIP learns only from text-image data, which lacks awareness of the sources’ audio traits. As an efficient solution, we learn a Cross-Modal Sound Source (CMSS) manifold as illustrated in Fig. 2 (b) to project the CLIP-CLAP embeddings to a joint semantic space where the local sound sources are disambiguated.
Finally, we attentively mix the CMSS embeddings together in Fig. 2 (c) to generate $\mathbf{a}$. This stage involves an attention-based Sound Source Remixer module.
Sound Source Perception
@@
Cross-Modal Sound Source Manifold
We contrastively learn the CMSS manifold from single-sound-source visual-audio pairs to project the perceived sound source semantics to a joint semantic space for disambiguation, as shown in Fig. 3 (a). The CMSS manifold naturally accommodates the multimodality of our perceptions due to the bridging of CLIP and CLAP.
@@
given a single-source visual-audio pair as $\left(\mathbf{V},\mathbf{A}\right)$ and its CLIP-CLAP embeddings as $\left(\mathbf{v},\mathbf{a}\right)$. $\mathbf{e}$ denotes the CMSS embedding. The projectors optimize a contrastive loss to attract visual-audio embeddings from the same sound-source pair and repel those from different sources. Following the symmetric contrastive guidance of CLAP (Elizalde et al. 2023), this objective can be formulated for a batch of $N$ pairs as:
@@
where $\lambda_{1}$ is a weight hyperparameter and $\mathcal{L}_{kl}$ is the summed K-L loss. During training, we model all three modules $\upsilon\left(\cdot\right)$, $\phi\left(\cdot\right)$, and $\chi\left(\cdot\right)$ with residually connected MLPs and alternatively optimize the projectors and generator with $\mathcal{L}_{\text{fold}}$.
@@
Method VGGSound Subjective Relevance Subjective Fidelity
[…]
S&H 10.929 79.221 5.941 1.886 - - - -
S&H-Text 5.087 31.461 8.690 2.646 2.358 0.845 2.460 0.748
Im2Wav 5.855 24.303 10.965 3.105 2.595 0.736 2.273 0.757
V2A-Mapper 0.946 5.516 11.521 3.164 3.063 1.024 2.693 1.210
SS2A(Ours) 1.150 6.716 12.947 3.793 4.080 0.527 4.098 0.459
Table 1: General image to audio tests. The VGGSound and subjective tests, without source annotation, generalize single-source and multi-source synthesis scenarios. The first and second places are bolded and underlined, respectively.
Sound Source Remixer
@@
We model $\psi\left(\cdot\right)$ with a stack of self-attention layers and learn it from visual-audio pairs in VGGSound. The visual sources are perceived from each video’s central frame following our aforementioned perception method. Each token sequence $\left\{\mathbf{x}_{m}\right\}$ is padded to a fixed length of $M=64$. To enhance generation diversity, a Classifier-free Guidance (Ho and Salimans 2021) is applied during training by randomly zeroing out tokens. We replace the classic attention with Efficient Attention (Shen et al. 2021) and detail this architecture in Supplementary Section 2.3. During inference, we set $\mathbf{v}=\mathbf{0}$ for sound source conditions from audio modality.